What is TFP and why is if important?

How do you calculate annual TFP growth rate?

How to calculate contributions of capital, labor, and TFP to total output?



You can make up numbers if it makes it easier to explain but please label what each value is
Answer #1

Total factor productivity refers to the efficiency of the factors of production in enhancing growth rate in the economy in addition to the output growth caused by the measired values of the inputs of production. It is important because it enables comparison of the efficiency of the factors of production in causing output growth across countries or regions or firms.

TFP growth rate is calculated as the ratio of total output produced and the total inputs used in its production.

Suppose the total output of bread in a bakery is of value $100

Suppose that labour worth $80 and fuel( capital) worth $10 were the only inputs used in the production of bread.

Therefore, the total factor productivity = Value of bread produced/ value of inputs used = 100 / 90 = 1.11

Now the additional output produced by the factors of production = 100 - 90 = 10

Total output = 100

value of labour = $80

value of capital = $10

Value of additional output due to TFP =$ 10

Share of labour in total output = 80 / 100 = 0.8

Share of capital in total output = 10 /100 = 0.1

Share of TFP in total output = 10/100 = 0.1